# LSM Trees in Adversarial Environments

### v1

@@ 6. Open Problems
[…]
General Adversary-resilience. In our work, we have shown a construction that provides adversary-resilience in an LSM store against $\mathbb{A}_{\text{BF}}$, the set of computationally bound Bloom Filter targeting adversaries. We leave answering the following follow-up questions as open problems:
@@
Learned Bloom Filters. We conjecture that replacing Bloom Filters used by an LSM store with Learned Bloom Filters (Bishop and Tirmazi, 2024) with a better false positive rate may lead to better performance. Adversary resilience for Learned Bloom Filters can be solved using the construction of (Bishop and Tirmazi, 2024). We leave experimental validation of this as an open problem. Similar research for Adaptive Bloom Filters (or Learned Adaptive Bloom Filters (Dai and Shrivastava, 2020)) will also be interesting.
Existence of an Ideal-World Simulator. In our simulator-based adversarial model (Appendix A), we have not proved the existence of an ideal-world simulator. We leave proving the existence of and providing a construction for an ideal-world simulator for LSM stores as an open problem. The constructions of an ideal-world simulator for probabilistic data structures by (Filić et al., 2022, 2024) may be a good starting point. However, their constructions require two properties in a data structure: function decomposability and reinsertion variance. Informally, reinsertion invariance requires the internal state of a probabilistic data structure to remain unchanged when the same key is reinserted. This does not necessarily apply to LSM stores.
@@ Appendix A Simulator-based Model
Appendix B Security Proofs
We first define the notion of secure Bloom Filters (Def. 4.1). We use known definitions (Filić et al., 2022; Tirmazi, 2025) for Bloom Filter (BF) constructions. A BF, $\Pi$ consists of two algorithms.
Game B.1 (Smash-Bloom).
Step 1:
Step 2:
Step 3:
Step 4:
Definition B.0.
[…]
Definition B.0.
Definition B.0.
Definition B.0.
$\leq\mathtt{negl}(n)$
Theorem B.5.
Proof.
  1. (1)
  2. (2)
[…]
We can now prove our main result regarding the security of LSM stores.
Theorem B.6.
Proof.
  1. (1)
    $I^{\prime}_{r}((k,v),\sigma)=C_{r}((F_{\kappa}(k),v),\sigma)$
  2. (2)
    $Q^{\prime}(x,\sigma)=Q(F_{\kappa}(x),\sigma)$.
Security Proof: We prove this by contradiction. Suppose there does exist an n.u. p.p.t. adversary $\mathcal{A}$ that can with the Smash-Lsm game with probability greater than $\varepsilon+\textrm{negl}(\lambda)$. Then by definition of oracle $\mathscr{O}_{Q}$, there exists a Bloom Filter $\Pi^{\prime}$ with the construction from Thm B.5 for which $\mathcal{A}$ can generate false positives with probability higher than $\varepsilon+\mathrm{negl}(\lambda)$. So $\mathcal{A}$ can also win the Smash-Bloom game with a probability higher than $\varepsilon+\mathrm{negl}(\lambda)$. This is a contradiction as Thm B.5 proves that no n.u p.p.t adversary can win the Smash-Bloom game against Bloom Filter constructions of type $\Pi^{\prime}$ with probability higher than $\varepsilon+\mathrm{negl}(\lambda)$.
Appendix C Adversarial Targets
There is a vast universe of adversarial targets we leave as open problems that emerge from our work. Empirical evidence showing performance degradation from these targets would help greatly in the design of future LSM stores.
Deleted Insertions. A direct follow-up for our attacks comes from the observation that Bloom Filters in LSM trees cannot perform deletions. This means that if an adversary $\mathcal{A}$ sabotages Bloom Filters via well-chosen insertions and then deletes the keys it inserted, the false positive rate of the Bloom Filters remains high unless the Bloom Filter is reconstructed from existing keys.
Range Queries. Range queries (Dayan et al., 2017; Kaushik and Sarkar, 2024) look for a larger number of keys in multiple levels of an LSM tree compared to point queries. Therefore, they are potentially more vulnerable to adversarial workloads. Adversary $\mathcal{A}$ can insert data in a manner that forces range queries to access an excessive number of runs, increasing read latency. Studying adversarial range query complexity and designing more efficient prefetching and merging strategies could help mitigate these attacks.
Merge Policy. There are different merge policies within LSM stores including leveling and tiering (Dayan et al., 2017). Prior work (Huynh et al., 2022) has shown that leveling is more robust to uncertain (but not explicitly adversarial) workloads than tiering. We conducted our evaluations on LevelDB and RocksDB, both of which use leveling. We might potentially see higher performance degradation on LSM stores that rely on tiering.

### v2

@@ 7. Open Problems
We leave the community with open problems in four categories that emerge from our work.
7.1. Adversarial Targets.
The experiments in our work focus on adversaries that target Bloom Filters in LSM stores via well-chosen insertions. We conduct our experiments on, LevelDB and RocksDB, both of which use a merge policy called leveling (Dayan et al., 2017; Huynh et al., 2022). There is a vast universe of adversarial targets we leave as open problems that emerge from our work. Empirical evidence showing performance degradation from these targets would help greatly in the design of future LSM stores.
Deleted Insertions
A direct follow-up for our attacks comes from the observation that Bloom Filters in LSM trees cannot perform deletions. This means that if an adversary $\mathcal{A}$ sabotages Bloom Filters via well-chosen insertions and then deletes the keys it inserted, the false positive rate of the Bloom Filters remains high unless the Bloom Filter is reconstructed from existing keys.
Range Queries
Range queries (Dayan et al., 2017; Kaushik and Sarkar, 2024) look for a larger number of keys in multiple levels of an LSM tree compared to point queries. Therefore, they are potentially more vulnerable to adversarial workloads. Adversary $\mathcal{A}$ can insert data in a manner that forces range queries to access an excessive number of runs, increasing read latency. Studying adversarial range query complexity and designing more efficient prefetching and merging strategies could help mitigate these attacks.
Merge Policy
There are different merge policies within LSM stores including leveling and tiering (Dayan et al., 2017). Prior work (Huynh et al., 2022) has shown that leveling is more robust to uncertain (but not explicitly adversarial) workloads than tiering. We conducted our evaluations on LevelDB and RocksDB, both of which use leveling. We might potentially see higher performance degradation on LSM stores that rely on tiering.
7.2. General Adversary-resilience
In our work, we have shown a construction that provides adversary-resilience in an LSM store against $\mathbb{A}_{\text{BF}}$, the set of computationally bound Bloom Filter targeting adversaries. We leave answering the following follow-up questions as open problems:
@@ 7.2. General Adversary-resilience
7.3. Learned Bloom Filters.
We conjecture that replacing Bloom Filters used by an LSM store with Learned Bloom Filters (Bishop and Tirmazi, 2024) with a better false positive rate may lead to better performance. Adversary resilience for Learned Bloom Filters can be solved using the construction of (Bishop and Tirmazi, 2024). We leave experimental validation of this as an open problem. Similar research for Adaptive Bloom Filters (or Learned Adaptive Bloom Filters (Dai and Shrivastava, 2020)) will also be interesting. A helpful starting point for this work is (Bishop and Tirmazi, 2024) who provide an adversary-resilient construction of the Learned Bloom Filter, called the Downtown Bodega Filter, in the same game-based setting as (Naor and Eylon, 2019).
7.4. Existence of an Ideal-World Simulator
In our simulator-based adversarial model (Appendix A), we have not proved the existence of an ideal-world simulator. We leave proving the existence of and providing a construction for an ideal-world simulator for LSM stores as an open problem. The constructions of an ideal-world simulator for probabilistic data structures by (Filić et al., 2022, 2024) may be a good starting point. However, their constructions require two properties in a data structure: function decomposability and reinsertion variance. Informally, reinsertion invariance requires the internal state of a probabilistic data structure to remain unchanged when the same key is reinserted. This does not necessarily apply to LSM stores.